[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Wed Apr 26, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Wed Apr 26 18:54:51 UTC 2017


TAFB Shift Log for Wed Apr 26 2017 14:54:45 GMT-0400 (EDT)

Shift: Day Shift

Pacific Forecaster: S Stripling

Atlantic Forecaster: C McElroy

Surface Analyst: A Latto

Technical Issues Summary:

GFE/NCP issues:

N-AWIPS issues: Starfish froze on CM upon reloading an spf. Does not have a  
heavy load of spf's on that box today. Hard reboot initiated.

Other issues:

Miscellaneous issues: Recieved a phone call from Fran at OPC, forwarding an  
urgent request for assistance in a Bahamas SAR. Told Fran to forward me  
info via email and have been since working to investigate, readying  
forwarded docs from NWS HQ, providing spot forecast, and re-runing HYSPLIT  
from initial time of incidence, which was Fri night, April 21. See email  
thread
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170426/eea6c2a1/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list